What mass of NaCl is needed to produce 500 mL of a 2.0 mol/L solution?
1
Expert's answer
2013-01-17T09:56:34-0500
1)Convert mL to L:
500mL=0.5L
2) find number of moles NaCl in our solution:
n(NaCl)=C(NaCl)*V(NaCl)=2.0mol/L * 0.5L=1.0 mol
3) find mass of NaCl in 1 mol:
m(NaCl)=n(NaCl)*Mr(NaCl)=1.0mol*58.44 g/mol=58.4 g.

answer:
we need 58.4 g of NaCl.